Solve by taking square roots:
2(x-3)^2=72

Answer by ewatrrr(24785)

 
Hi
Solve by taking square roots:
2(x-3)^2=72
(x-3)^2= 36
(x-3) = ± 6
x = 3 ± 6
Two solutions: x = 9 and x = -3